### v2

# Thermodynamic uncertainty relations in superconducting junctions

Abstract
It has been recently predicted that the thermodynamic uncertainty relation (TUR), which expresses a trade-off between the amount of dissipation and the absence of fluctuations, can be violated in the context of charge transport in normal quantum conductors. However, the required conditions are difficult to meet and no experimental violation has been yet reported in mesoscopic quantum transport experiments. Here, we predict that TUR can be largely violated in superconducting contacts without introducing an energy dependency in the normal transmission coefficient. This violation originates from the coexistence of tunneling processes involving different numbers of transmitted charges, namely quasiparticle tunneling and Andreev reflections. The superconducting junctions invoked here have been fabricated by many groups and constitute an ideal platform to demonstrate TUR’s breaking in charge transport experiments.
Introduction.— The implementation of superconducting materials has become a key ingredient to design better and more functional quantum thermal machines. They have been proposed to be efficient for cooling performances [nahum1994electronic, leivo1996efficient, PhysRevB.98.241414, PhysRevB.107.245412, PhysRevResearch.5.043041], as well as for pumping or generating power [PhysRevB.106.115419, PhysRevB.103.235434]. Besides, superconductors can exhibit remarkable thermoelectric properties in the nonlinear transport regime when electron-hole symmetry is broken [PhysRevLett.124.106801, germanese2022bipolar, PhysRevResearch.6.L012022, PhysRevB.106.245413].
Another advantageous aspect when dealing with superconducting elements is that they overcome the fundamental bound imposed by the thermodynamic uncertainty relation (TUR) [Barato2015, Todd2016, Horowitz2020, Pietz2016, Pietzonka16b, Pietz2018, Timp2019, PhysRevB.98.155438, PhysRevResearch.5.043041]. The TUR establishes a trade-off relation between uncertainty of a current traversing the system and the associated entropy production or dissipation. The TUR is a nonequilibrium relation derived from the Markovian evolution of a system and sets a more restrictive bound for the entropy production in terms of the noise to current ratio. In the case of charge transport, TUR dictates $S/I^{2}\geq 2k_{\rm B}/\dot{\Sigma}$ with $k_{\rm B}$ the Boltzmann constant, $\dot{\Sigma}$ the entropy production, $I$ the average charge current and SS the corresponding current fluctuations.
TUR violations have been predicted in quantum systems attributed to either quantum coherence or the breakdown of the local detailed balance [Andrieux_2007, PhysRevLett.120.090601, PhysRevX.11.021013, PhysRevResearch.5.043041, PhysRevResearch.5.013038, PhysRevResearch.5.023155, PhysRevB.28.1655]. Large deviations from the TUR are desirable for the quantum-enhanced performance of the engine operation, i.e., less dissipative and more stable machines [PhysRevB.98.085425, PhysRevB.98.155438, PhysRevLett.120.090601, PhysRevE.99.062141]. Besides, quite recently there has been a notable effort towards finding quantum systems exhibiting a prominent breakdown of the TUR [PhysRevResearch.5.043041, PhysRevResearch.5.023155, PhysRevB.104.045424, PhysRevResearch.5.013038, PhysRevResearch.5.023155, PhysRevB.108.115422]. Thus, for instance, a chain of quantum dots with a rectangular transmission was proposed to violate the standard TURs [PhysRevB.104.045424]. Additionally, moderate deviations in the TUR relation have been predicted in hybrid normal-superconductor quantum dots in the weak tunnel coupling regime [PhysRevResearch.5.043041]. However, all these proposals require stringent conditions and all attempts to observe a violation of TUR in quantum conductors have failed thus far, see e.g. Ref. [PhysRevB.101.195423].
Here, we predict that TUR can be largely violated in few-channel quantum point contacts involving superconducting leads, which have already been realized e.g. with the help of superconducting atomic-size contacts. We show that the strong violations of TUR in these junctions are due to the coexistence of different transport processes, namely single-quasiparticle tunneling (QP) and (multiple) Andreev reflections (AR). We find that the largest TUR violations occur for highly transmissive contacts and no energy dependence of the normal state transmission is required, contrary to what happens in normal quantum point contacts [PhysRevB.98.085425, PhysRevB.104.045424].
Quantifying TUR deviations.— In what follows we present a general discussion to explain what makes superconductors special in the context of TUR. For this purpose, we consider a generic isothermal two-terminal quantum device that may contain normal metals (N) and/or superconductors (S). If we apply a bias voltage $V$, entropy production due to the Joule heating is given by $\dot{\Sigma}=IV/T$, where $I$ is the charge current and $T$ the temperature of the electrodes. Deviations for the TUR, i.e., $\frac{S}{I^{2}}\dot{\Sigma}\geq 2k_{B}$ with $S$ being the current fluctuations, can be quantified by the TUR-breaking coefficient
$\mathcal{F}\equiv F^{*}-\frac{2k_{\rm B}T}{V},$ (1)
where $F^{*}=S/I$ is the Fano factor and we have assumed that the electron charge is unity. The coefficient $\mathcal{F}$ is negative for broken TUR and positive (or zero) for unbroken TUR, where $\mathcal{F}=0$ corresponds to the classical limit. Below, we shall present a quantitative analysis of $\mathcal{F}$ valid for arbitrary range of parameters, but to get an insight into the conditions required to break TUR, we now follow Ref. [PhysRevB.98.085425] and focus on the low-bias regime in which it is possible to make analytical progress. With this idea in mind, we make a Taylor expansion in the bias of the current and noise
$I$ $=$ $G_{1}V+\frac{1}{2!}G_{2}V^{2}+\frac{1}{3!}G_{3}V^{3}+\mathcal{O}(V^{4}),$ (2)
$S$ $=$ $S_{0}+S_{1}V+\frac{1}{2!}S_{2}V^{2}+\mathcal{O}(V^{3}).\quad\quad$ (3)
Here, $G_{1},G_{2},G_{3}$ are linear and nonlinear conductances and $S_{0},S_{1},S_{2}$ are the equilibrium and nonequilibrium noise terms. Using the nonlinear fluctuation relations [PhysRevLett.101.136805, PhysRevB.72.235328, Andrieux_2007, PhysRevLett.108.246603] together with the Johnson-Nyquist relation for the equilibrium noise (i.e., $S_{0}=2k_{\rm B}TG_{1}$), we obtain $\mathcal{F}=Vk_{\rm B}T/(G_{1})C_{\rm neq}+\mathcal{O}(V^{2})$ where the coefficient [PhysRevB.98.155438]
$C_{\text{neq}}\equiv\frac{3S_{2}-2k_{\rm B}TG_{3}}{6k_{\rm B}T}$ (4)
quantifies TUR deviations. Namely when $C_{\text{neq}}<0$ (in the limit of small voltages), TUR is violated.
The stochastic nature of charge transfer in mesoscopic junctions allows us to express the current in an arbitrary mesoscopic junction as follows [Levitov1996, Klich2003]
$I=\int\frac{dE}{h}\left(\sum_{n}np_{n}\right),$ (5)
where $p_{n}=p_{n}(E,V,T)$ are charge-resolved probabilities of a transport process transferring $n$ charges across the junction that depends on the energy $E$, voltage $V$ and temperature $T$. The corresponding current fluctuations are given by [Levitov1996, Klich2003]
$S=\int\frac{dE}{h}\left[\sum_{n}n^{2}p_{n}-\left(\sum_{n}np_{n}\right)^{2}\right].$ (6)
These formulas for $I$ and SS hold for junctions containing normal metals and/or superconductors. For normal contacts, the transport is solely due to single-electron tunneling ($|n|=1$), while with S electrodes multiple electron ($|n|>1$) tunneling processes can coexist. This is what makes superconductors special, as we are about to show.
To get further insight, we expand the charge-resolved probabilities in the applied voltage
$p_{n}=p_{n}^{(0)}+p_{n}^{(1)}V+p_{n}^{(2)}V^{2}/2+p_{n}^{(3)}V^{3}/6+\mathcal{O}(V^{4}).$ (7)
By using the detailed balance condition $p_{-n}=p_{n}e^{-n\beta V}$ [PhysRevB.72.235328] where $\beta=1/(k_{\rm B}T)$ and inserting the expanded probabilities [see Eq. (7)] in the current and noise formula [see Eqs. (5), and (6)] we obtain the following result for the coefficient $C_{\rm neq}$ in Eq. (4)
$\frac{6C_{\rm neq}}{\beta^{3}}=\int\frac{dE}{h}\Bigg($ $\sum_{n>0}n^{4}p_{n}^{(0)}(1-6p_{n}^{(0)})$
$-\sum_{m>n>0}12n^{2}m^{2}p_{n}^{(0)}p_{m}^{(0)}\Bigg),$ (8)
which notably only depends on the zeroth order approximation of the transmission probabilities $p_{n}^{(0)}$.
To exemplify Eq. (Thermodynamic uncertainty relations in superconducting junctions) in the context of superconducting junctions, we now consider a single-channel NS contact. The superconducting lead is assumed to be a conventional BCS superconductor with a temperature-dependent gap $\Delta\equiv\Delta(T)$. In this case, there are two transport processes as illustrated in Fig. 1(a,b). For zero temperature, QPs can only be transferred for energies above the gap ($V\geq\Delta$) and transfer one charge which is shown in panel Fig. 1(a). In the subgap region ($V\leq\Delta$) transport can take place via Andreev reflection where two charges are transferred in one instance, see panel Fig. 1(b). For non-zero temperatures, QP transfer can also occur for in-gap voltages because of the thermal broadening of the corresponding Fermi functions. The two processes are associated with different transmission probabilities, namely $p_{\pm 1}$ corresponds to QP tunneling and $p_{\pm 2}$ to AR and all other probabilities are zero [Belzig2003]. With this information at hand, the coefficient $C_{\rm neq}$ reads:
$\frac{6C_{\rm neq}^{\rm NS}}{\beta^{3}}{=}\int\frac{dE}{h}\left[p_{1}^{(0)}(1{-}6p_{1}^{(0)}){+}2^{4}p_{2}^{(0)}(1{-}6p_{2}^{(0)}){-}48p_{1}^{(0)}p_{2}^{(0)}\right].$ (9)
This expression reduces to the result in Ref. [PhysRevB.98.155438] for normal contacts ($p_{2}^{(0)}=0$). The new contributions in the NS case, second and third terms, are related to the AR and the coexistence of the two processes, respectively. The sign of $C_{\rm neq}^{\rm NS}$ depends on the precise energy dependence of $p^{(0)}_{1,2}$. One can gain some intuition by analyzing an NS contact with an energy-independent normal transmission coefficient. In this case the $p^{(0)}_{1,2}$ coefficients can be computed e.g. with the well-known BTK model [PhysRevB.25.4515]. As we show in Ref. [SM], such an analysis suggests that two requirements must be simultaneously met to significantly break TUR: (i) coexistence of the two processes at a given bias and temperature and (ii) high transmissions. The coexistence of the two processes is responsible for a large negative contribution of the cross term $48p_{1}^{(0)}p_{2}^{(0)}$ in Eq. (9), while high transmissions are necessary to have high probabilities and make the terms $n^{4}p_{n}^{(0)}(1-6p_{n}^{(0)})$ negative. The dependence of the zero-bias probabilities on transmission and temperature, as well as a discussion on the role of the different terms in Eq. (9), can be found in Ref. [SM]. This intuition will be further confirmed below by our numerically exact results at arbitrary bias and we shall use those results to test the validity of Eq. (9).
In the case of SS contacts, new transport processes in the form of multiple Andreev reflections (MAR) can take place. Their corresponding probabilities $p_{\pm n}$ for $n>2$ contribute to the current and noise (see Ref. [Cuevas2003, Cuevas2004]) and therefore appear in Eq. (Thermodynamic uncertainty relations in superconducting junctions). The lowest order transport processes in an SS junction are depicted in Fig. 1(c-f) and correspond to QP ($p_{\pm 1}$), AR ($p_{\pm 2}$), MAR ($p_{\pm 3}$) and 4th order MAR ($p_{\pm 4}$) transferring one, two, three and four electrons respectively.
Figure 1: (a) Single-QP tunneling in which a QP tunnels into the empty density of states of the superconductor transferring one electron. The applied voltage has to exceed the SC gap, $V\geq\Delta$. (b) AR in which an electron is reflected as a hole inside the superconducting gap transferring two electrons. The voltage is arbitrary, $V\geq 0$. (c) QP tunneling between two SCs at voltages above $2\Delta$. (d) AR from a voltage biased superconductor to another superconductor transferring two electrons. The voltage has to exceed the SC gap, $V\geq\Delta$. (e) First MAR, where an incoming electron is reflected as a hole and that hole is retro-reflected as another electron, transferring a total of three electrons. The voltage has to exceed $V\geq 2\Delta/3$. (f) Higher order MAR involves a process that transfers four electrons with onset voltage $V\geq 2\Delta/4$. The voltage thresholds correspond to the case of zero temperature.
QPC of constant transmission.— In the following we focus on single-channel NS and SS contacts characterized by an energy-independent normal state transmission coefficient $\tau$. These junctions have been realized with the help of superconducting atomic-size contacts in which both the current and the noise have been thoroughly studied [Scheer1997, Scheer1998, Cuevas1998, Cron2001, Senkpiel2020, Senkpiel2020a]. We make use of the framework of full counting statistics to obtain current, noise and charge resolved probabilities using Ref. [Ohnmacht2023] to compute the full result for the TUR-breaking coefficient in Eq. (1).
Figure 2: (a) Critical transmission $\tau_{c}$ for breaking TUR for a NS junction as a function of the bias voltage $V/\Delta$ and the temperature $k_{\rm B}T/\Delta$. For areas with no color, TUR is not broken. (b) The minimal TUR breaking coefficient $\mathcal{F}_{\rm min}$ as a function of the dimensionless temperature $k_{\rm B}T/\Delta$ and the transmission $\tau$ for a NS junction. A negative coefficient indicates broken TUR. The black solid line indicates the phase boundary. The dotted line indicates the approximate phase boundary from the coefficient $C_{\rm neq}$ [see Eq. (9)]. The inset shows a zoom-in into the region where TUR is broken. (c) Same as in (a) for an SS junction. Dotted lines indicate special voltages of onsets of different transport processes, namely $V/\Delta=1,2/3,1/2$ for AR, MAR and 4th order MAR. Grey area shows a parameter realm that was not analyzed. (d) Same as in (b) for an SS junction with the phase boundary as a solid line. The inset shows a zoom-in into the region where TUR is broken.
Our purpose is to investigate whether TUR can be broken in a NS and SS system and portray the “phase” diagram of the system. We assign the phase of unbroken TUR if $\mathcal{F}\geq 0$ and the second phase of broken TUR if $\mathcal{F}<0$. To illustrate the results, we fix the voltage and temperature and vary the transmission $\tau$ of the QPC. We define $\tau_{\rm c}$ as the critical transmission for which the lower bound of TUR inequality is satisfied ($\mathcal{F}=0$). It holds that for any $\tau<\tau_{\rm c}$, TUR is fulfilled whereas it is violated for $\tau>\tau_{\rm c}$. In Fig. 2(a) we show $\tau_{\rm c}$ as a function of the dimensionless temperature $k_{\rm B}T/\Delta$ and voltage $V/\Delta$ for the NS case. Notice that the TUR only breaks at high transmissions ($\tau_{c}\approx 0.91$), as anticipated in the discussion of Eq. 9. Note also that $\tau_{\rm c}$ decreases upon increasing the temperature, which is due to the larger overlap between the two process induced by the finite temperature. We want to stress that no energy dependence of the transmission is needed for breaking TUR. The source of nonlinearity of the transport characteristics ($I$ and $S$) as a function of the bias that is required to violate TUR arises from the coexistence of QP tunneling and AR in addition to the non-constant density of states of the SC. This is in contrast to the results in Ref. [PhysRevB.98.155438], where an energy dependence of the transmission is needed. For low temperatures, TUR is only broken for $V\approx\Delta$ and only for $\tau\approx 1$. This indicates that the breaking of TUR is rooted in the coexistence of the QP and AR (for more details see Ref. [SM]).
To quantify the violation of TUR, we consider the voltage at which TUR is maximally broken and define the minimal TUR-breaking coefficient $\mathcal{F}_{\rm min}\equiv{\rm min}_{V}\mathcal{F}$, which is positive (or zero) for unbroken TUR and negative for broken TUR. In addition, the more negative $\mathcal{F}_{\rm min}$ the more TUR is broken. In Fig. 2(b), this coefficient is plotted as a function of the transmission and the temperature. The solid line indicates the phase boundary between the unbroken and broken TUR phase. Notice that the TUR is more broken the higher the transmission is. These findings can be understood with the help of the low-bias analysis above, see Eq. (9). The probabilities $p_{1,2}$ are analytical in the NS (see Ref. [Ohnmacht2023, SM]) and $C_{\rm neq}^{\rm NS}$ can be easily calculated. The phase boundary obtained with this approximation is shown as a dotted line in panel Fig. 2(b). It qualitatively reproduces the exact results, especially at high temperatures where the energy dependence of $p_{1,2}$ is mainly determined by the smearing of the Fermi functions (for more details see Ref. [SM]). On the other hand, the fact that high transmissions are required to have a stronger violation of TUR can be understood from the contribution $2^{4}p^{(0)}_{2}(1-6p^{(0)}_{2})$ in Eq. (9). This term scales quartically with the charge, which renders a stronger violation for highly probable AR processes coexisting with low probability QP processes. Furthermore, an important contribution to $C_{\rm neq}^{\rm NS}$ comes also from the cross term $48p_{1}^{(0)}p_{2}^{(0)}$.
For completeness, we also address in Ref. [SM] the case of an NS contact featuring an energy dependent transmission coefficient. In particular, we have studied the case in which the leads are coupled via a quantum dot. The transmission coefficient in this case corresponds to a Breit-Wigner resonance centered in the energy of the dot level $\epsilon_{0}$ with a tunneling broadening given by $\Gamma$, i.e., $\tau(E)=\Gamma^{2}/[(E-\epsilon_{0})^{2}+\Gamma^{2}]$. However, we find that this added energy dependence does not further break TUR, see Ref. [SM], which we attribute to the fact that the resonant level tends to reduce the high effective transmission at some energies that is needed to break TUR.
We turn now to the analysis of the SS case, where the presence of MAR processes, which are shown in Fig. 1(e-f), adds additional complexity. Again we refer to Ref. [Cuevas2003, Cuevas2004, Ohnmacht2023] for the current and noise calculation in the framework of the full counting statistics. The critical transmission $\tau_{\rm c}$ for the SS case is shown in Fig. 2(c). It is noticeable that the TUR is broken for generally smaller transmissions in comparison to the NS case [see panel (a)]. For higher temperatures, the critical transmission decreases because the processes overlap more, which benefits the TUR breaking. As compared to the NS case, the critical transmission depends more drastically on the voltage even for higher temperatures. In particular, the dependence of the voltage is more pronounced in the vicinity of voltages corresponding to onsets of MAR processes, namely $V=2\Delta/n$, with $n$ being an integer. Also notice that for low temperatures, the TUR violation concentrates on only one point, namely $V=T=0$. This can be explained by considering that in the limit $V,T\to 0$ for an SS junction with high transmission, all transmission probabilities up to infinite order contribute. In the case of perfect transmission, the current does not go to zero for $V,T\to 0$ and converges to a non-zero value [Cuevas2003, Cuevas2004], rendering the absolute limit of violation.
The minimal TUR breaking coefficient is displayed in Fig. 2(d), where it is seen that TUR is severely more broken than in the NS case, reaching absolute values of one order of magnitude higher than in the NS case. This enhanced breaking must originate from the MARs, which are exclusive to the SS. In particular, this large breaking results from highly probable, high-order MAR processes coexisting with lots of lower probability MARs of lower order, granting an immensely strong violation of the TUR. As the voltage range is restricted for $\mathcal{F}_{\rm min}$ in Fig. 2(d) due to computation time, $V\geq 0.1\Delta$, an even stronger violation is achievable for lower voltages.
Conclusions.– In this work, we have studied the violation of TUR in the coherent charge transport in junctions containing superconductors. We have shown that the presence of multiple tunneling processes help breaking TUR and that the degree of violation is clearly larger in the SS case. We have rationalized this by providing an analytical formula for the TUR-breaking coefficient for low voltages and show that it only depends on the zero-bias tunneling probabilities. We have also shown that for junctions containing superconductors, no energy dependence of the normal transmission coefficient is required to break TUR, which is due to the intrinsic energy dependence of the probability of the different superconducting tunneling processes. This is a key advantage that makes it much more feasible to observe the violation of TUR in superconducting junctions. So, in summary, our work shows that superconducting junctions of simple quantum point contacts are promising candidates for the engineering of quantum thermal machines that are immune against fluctuations in a low dissipation scenario. Our contribution also paves the way for the experimental observation of the TUR breaking in the context of quantum electron transport, something that continues to be very elusive.
D.C.O. and W.B. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG; German Research Foundation) via SFB 1432 (Project No. 425217212). J.C.C. acknowledges support from the Spanish Ministry of Science and Innovation under Grant No. PID2020-114880GB-I00 and through the “María de Maeztu” Programme for Units of Excellence in R&D (CEX2023-001316-M) and thanks the DFG and SFB 1432 for sponsoring his stay at the University of Konstanz as a Mercator Fellow. R.L. acknowledges support by the Spanish State Research Agency (MCIN/AEI/10.13039/501100011033) and FEDER (UE) under Grant No. PID2020-117347GB-I00 and María de Maeztu Project No. CEX2021-001164-M and Grant No. PDR2020/12 sponsored by the Comunitat Autonoma de les Illes Balears through the Direcció General de Política Universitaria i Recerca with funds from the Tourist Stay Tax Law ITS 2017-006. K.W.K. acknowledges the support by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2020R1A5A1016518).